OpenStudy (anonymous):

given the acceleration, initial velocity, and initial position of a body moving along a coordinate line at time t, find the body's position at time t a=12, v(0)=8, s(0)=1

OpenStudy (anonymous):

just integrate starting with the acceleration: \(\large \int 12 dt= \) ???? this will give you the velocity. use the intial velocity to find the constant of integration. then integrate this velocity to find the position function. use the initial position to determine the constant of integration.

OpenStudy (anonymous):

integrate 12dt = ? sorry I don't know

OpenStudy (anonymous):

you don't know how to integrate?

OpenStudy (anonymous):

\[a = \frac{v}{t} \text{ and } v = \frac{s}{t}\] Knowing that, you can solve for \(t\) without integrating.

OpenStudy (anonymous):

thanks